OpenStudy (anonymous):

Find all solutions to the equation. (sin x)(cos x) = 0 a.{npi|n=0,1,2,...} b.{(pi/2)+npi,npi|n=0,1,2,...} c.{(pi/2)+2npi|n=0,1,2,...} d.{(pi/2)+npi|n=0,1,2,...}

OpenStudy (anonymous):

can anyone please help? ;(

OpenStudy (anonymous):

\(\sin(x)=0\) means \(x=0,\pi,2\pi,...\)

OpenStudy (anonymous):

\(\cos(x)=0\) means \(x=\frac{\pi}{2},\frac{3\pi}{2},\frac{5\pi}{2},...\)

OpenStudy (anonymous):

@satellite73 so it will be an a?

OpenStudy (anonymous):

taken together, it would be \(n\pi\) and \(\frac{\pi}{2}+\pi\)

OpenStudy (anonymous):

no, go with \(b\)

OpenStudy (anonymous):

\(\{\frac{\pi}{2}+n\pi, n\pi |n\in \mathbb{Z}\}\)

OpenStudy (anonymous):

@satellite73 thankyou!!:):)

OpenStudy (anonymous):

or if you like \(n=0,1,2,...\) but that is not really correct because \(n\) could be negative

OpenStudy (anonymous):

yw